# Towards Comprehensive and Efficient Post Safety Alignment of Large Language Models via Safety Patching

### v1

@@ F.3 Comparison of Training and Inference Costs
As shown in Table 9, unlearning-based methods (GA [Yao et al., 2023], GA + Mismatch [Yao et al., 2023], NPO [Zhang et al., 2024a]) require additional training overhead to maintain the model’s overall utility. This involves the inclusion of extra general data $D_{g}$ and the need for more extensive GPU memory usage. Specifically, $D_{g}$ is used to reduce the divergence between the output on $x^{g}$ from the target LLM and the original LLM. Formally, they introduce a KL divergence loss in addition to the gradient ascent loss function in Equation (1):
@@
Thus, during the training process, the original LLM also needs to be loaded onto the GPU for the computation in the above equation, which requires additional GPU memory usage.
@@
In Table 9, we showcase the inference speeds of various methods, measured using the average token generation time ratio (ATGR) metric [Xu et al., 2024]:
@@
It is evident that these decoding-based methods reduce the model’s inference speed due to additional operations on token probabilities during the decoding process. Although it serves as a method to enhance the safety of the base model during the decoding phase and has relatively low additional inference overhead, it still requires training a safety expert model using additional safety data to guide the probabilities of generating different tokens during the decoding process of the base model. Conversely, our SafePatching does not change the backbone’s architecture or inference process, resulting in no additional overhead and no need for further adjustments during model deployment.
@@
Figure 6: The performance in terms of safety, over-safety, and utility of our SafePatching under different parameter retention rates. The backbone is LLaMA-2-Chat-7B.
F.4 Impact of Overall Retention Rate
@@
In Figure 6, we show the model’s performance across our three PSA goals, safety enhancement, over-safety mitigation, and utility preservation, at different level of overall retention rates $p$. From this, we can derive four key insights:
Utility Preservation: When the overall retention rate increases from 0 to 50, the utility of the resulting PSA model consistently surpasses that of a model without the random retention mechanism (“w/o Rand. Retention” in Table 2). This underscores the importance of this mechanism in preserving the model’s utility.
Balancing Utility and Specific Patch Characteristics: The overall retention rate can balance the utility preservation of the resulting model and the retention of each patch’s characteristics. When the overall retention rate is 0, the parameters involved in patching only include the difference set of the most important parameter regions of the safety patch and the over-safety patch. As a result, with fewer patch parameters participating in the patching process, the resulting model achieves the best performance on the utility benchmarks (the best AVG.). However, there is no significant improvement in safety performance or in mitigating over-safety.
Increased Conflict with Higher Retention Rates: Excessive overall retention rates worsen the conflict between safety and over-safety. As the overall retention rate increases, more parameters beyond the importance difference set participate in the patching process, increasing the conflict between safety and over-safety.
Inverse Relationship with MT-Bench Performance: The resulting PSA model’s performance on MT-Bench inversely correlates with its over-safety trend. As the model’s over-safety issues are mitigated, it tends to provide more direct answers to user questions on MT-Bench, making it more helpful. Specific examples of this can be seen in Figure 9.
Figure 7: Visualization of the difference set of the most important parameters from the safety patch. The backbone is LLaMA-2-Chat-7B.
Figure 8: Visualization of the difference set of the most important parameters from the over-safety patch. The backbone is LLaMA-2-Chat-7B.
F.5 Visualization of Patching on the Difference Set
We show the difference set between the most important parameter regions for safety patch and over-safety patch in Figures 7 and 8, respectively. Our analysis reveals two main findings.
Firstly, the important parameters related to over-safety are mostly concentrated in the lower transformer layers of the backbone model, whereas those related to safety are relatively more evenly distributed, primarily in the middle layers of the model. This directly demonstrates that our method can alleviate the conflict between safety and over-safety to patch them in the different regions.
Secondly, the important parameters for safety are predominantly located in the FFN layers, while those for over-safety are more concentrated in the attention layers. This supports the findings of Shi et al. [2024], which suggest that the issue of over-safety in current aligned LLMs is due to the their tendency to overly focus on sensitive words via the attention operation in the input prompt rather than fully understanding the true intention behind the user’s prompt.
Appendix G Case Study
G.1 Case Study for Safety Enhancement
@@
G.2 Case Study for Over-Safety Mitigation
@@
G.3 Case Study for Utility Preservation (Increasing)
@@
Firstly, the primary reason SafePatching enhances model performance on MT-bench is that it effectively reduces the issue of over-safety in the backbone model, allowing it to directly address the user’s questions. As illustrated in Figure 9, when tasked with named entity recognition (NER), the original model, due to over-safety concerns, would first critique the phrasing of the user’s query (the input sentence of NER). This response is irrelevant to the task at hand, as highlighted by GPT-4. Our SafePatching, however, successfully addresses this issue, directly meeting the user’s harmless request, and is thus more favored by GPT-4.
@@
Appendix H Broader Impact
Our SafePatching have the potential to impact society in both positive and negative ways.
Positive Societal Impacts:
  • Enhanced Safety and Trust: By addressing the safety challenges inherent in current LLMs, SafePatching could make these models safer to use, thereby increasing public trust in AI. This would encourage wider adoption of LLMs in various applications, from customer service to educational tools, knowing that the risk of harmful or inappropriate responses is minimized.
  • Scalability and Efficiency in Post Safety Alignment: SafePatching’s efficiency in post safety alignment makes it easier to implement safety measures across different models and scenarios, potentially leading to a standardized approach to AI safety. This can streamline regulatory compliance and foster innovation while ensuring ethical AI practices.
Negative Societal Impacts:
  • Potential for Misuse: As with any powerful technology, there is a risk that enhanced LLMs could be misused. Even with improved safety mechanisms, malicious actors might find ways to exploit these models for disinformation, fraud, or other harmful purposes. Ensuring robust safeguards and monitoring systems will be crucial.
  • Privacy Issues:As safer LLMs would interact more seamlessly with users, the collection and handling of data become critical. Ensuring that user data is protected and that privacy is maintained is essential to prevent breaches and misuse of personal information.
Figure 9: Case study of the generated responses from the orignal backbone model and that from our SafePatching. We also provide the rating and rational from the GPT-4 evaluator.

### v2

@@ G.2 Comparison of Training and Inference Costs
As shown in Table 12, unlearning-based methods (GA (Yao et al., 2023), GA + Mismatch (Yao et al., 2023), NPO (Zhang et al., 2024a)) require additional training overhead to maintain the model’s overall utility. This involves the inclusion of extra general data $D_{g}$ and the need for more extensive GPU memory usage. Specifically, $D_{g}$ is used to reduce the divergence between the output on $x^{g}$ from the target LLM and the original LLM. Formally, they introduce a KL divergence loss in addition to the gradient ascent loss function in Equation 1:
@@
Thus, during the training process, the original LLM also needs to be loaded onto the GPU for the computation in the above equation, which requires additional GPU memory usage. For $\text{RESTA}_{d}$, the process includes fine-tuning the backbone on the English Alpaca dataset, followed by fine-tuning on $D_{h}$ to derive safety vectors. This approach also introduces additional training data.
@@
In Table 12, we showcase the inference speeds of various methods, measured using the average token generation time ratio (ATGR) metric (Xu et al., 2024):
@@
It is evident that these decoding-based methods reduce the model’s inference speed due to additional operations on token probabilities during the decoding process. Although SafeDecoding serves to enhance the safety of the base model during the decoding phase and has relatively low additional inference overhead, it still requires training a safety expert model using additional safety data to guide the probabilities of generating different tokens during the decoding process of the base model. Conversely, our SafePatching does not change the backbone’s architecture or inference process, resulting in no additional overhead and no need for further adjustments during model deployment.
@@
Figure 4: The performance on safety, over-safety, and utility of SafePatching under different parameter retention rates. The backbone is LLaMA-2-7B-Chat.
Figure 5: The performance in terms of safety, over-safety, and utility of our SafePatching under different top rate and scale weight. The x-axis represents the setting of $(a,b)$ and $(\alpha,\beta)$. The settings adopted in our main experiments are in bold and the backbone is LLaMA-2-7B-Chat.
Figure 6: The performance in terms of safety, over-safety, and utility of our SafePatching under different top rate and scale weight. The x-axis represents the setting of $(a,b)$ and $(\alpha,\beta)$. The settings adopted in our main experiments are in bold and the backbone is LLaMA-3-8B-Instruct.
Figure 7: The performance in terms of safety, over-safety, and utility of our SafePatching under different top rate and scale weight. The x-axis represents the setting of $(a,b)$ and $(\alpha,\beta)$. The settings adopted in our main experiments are in bold and the backbone is Gemma-1.1-7B-it. When the performance result is 0, it indicates that we are unable to obtain meaningful outcomes, which means the overall performance of the model has been severely compromised.
Figure 8: The performance in terms of safety, over-safety, and utility of our SafePatching under different top rate and scale weight. The x-axis represents the setting of $(a,b)$ and $(\alpha,\beta)$. The settings adopted in our main experiments are in bold and the backbone is Mistral-7B-Instruct-v0.1. When the performance result is 0, it indicates that we are unable to obtain meaningful outcomes, which means the overall performance of the model has been severely compromised.
G.3 Hyperparameter Analysis
We conduct detailed analysis on the robustness and stability for hyper-parameters in SafePatching, involving overall retention rate $p$ (Equation 5), top rate $a$ and $b$ (Equation 3.2), and scale weight $\alpha$ and $\beta$ (Equation 8).
Analysis on overall retention rates $p$
@@ G.3 Hyperparameter Analysis
In Figure 4, we show the model’s performance across our three PSA goals, safety enhancement, over-safety mitigation, and utility preservation, at different level of overall retention rates $p$. From this, we can derive two key findings:
  • When $p$ is high, most parameters of the Safety Patch and Over-Safety Patch conflict with each other. If they are directly merged, their effects may cancel each other out, causing the utility to approach that of the original model.
  • As $p$ decreases, the proportion of important parameters increases, and the merged model tends to exhibit more of the Patch characteristics, leading to a slight decrease in utility.
Through our experimental results, we find that $p=30$ yields the most balanced outcome. Thus, we fix this hyper-parameter for other backbones to reduce the complexity of hyper-parameter tuning. The following results also demonstrated the suitability of $p=30$.
Analysis on top rate $a$ and $b$
Results on LLaMA-2-7B-Chat, LLaMA-3-8B-Instruct, Gemma-1.1-7B-it and Mistral-7B-Instruct-v0.1 are displayed in the left part of Figure 5, Figure 6, Figure 7 and Figure 8, respectively. We keep the other three hyper-parameters $p$, $\alpha$ and $\beta$ consistent with the main experiment settings, only changing the values of $a$ and $b$. The stable results from the 9 groups demonstrate the robustness of these two hyper-parameters, which we set to 3 and 2 in all experimental settings and across different backbones.
Analysis on scale weight $\alpha$ and $\beta$
They are the only hyper-parameters that need to be adjusted to further balance safety and over-safety patches. This process involves merely assignment operations in storage space, making it very quick and efficient without involving any GPU computations. Results on LLaMA-2-7B-Chat, LLaMA-3-8B-Instruct, Gemma-1.1-7B-it and Mistral-7B-Instruct-v0.1 are displayed in the right part of Figure 5, Figure 6, Figure 7 and Figure 8, respectively.
To sum up, we can draw two conclusions from the above hyper-parameter analysis:
  • Random retention rate $p$ and the top rate $a$ and $b$ are robust across different backbones. And top rate also exhibits robustness within each single backbone.
  • The only model-specific hyper-parameter we need to adjust is the scale weight $\alpha$ and $\beta$, and: (1) Larger $\alpha$ and $\beta$ can negatively impact overall model performance and lead to meaningless responses (denoted as 0 in the above Figures), potentially leading to a loss of utility (especially on Mistral and Gemma). (2) Within an appropriate value range, $\alpha$ and $\beta$ further balance the conflict between safety and over-safety: the larger $\alpha$, the more pronounced the effect of the Safety Patch, enhancing safety performance but increasing over-safety. Conversely, a larger $\beta$ has the opposite effect.
G.4 Continual Post-Safety Alignment
We evaluate SafePatching under a more challenging and practical continual PSA setting.
Experimental Setting for Continual Safety Alignment
We select three safety categories from Beavertails to simulate user safety alignment requirements at three consecutive time steps. Specifically, these three safety categories are “drug_abuse,weapons,banned_substance”, “animal_abuse”, and “misinformation_regarding_ethics,laws_and_safety”, each with 1,000 harmful data points for training and 100 for testing. At each time step, we train SafePatching for 1 epoch with gradient ascent and 3 epochs with gradient decent, with the total batchsize of 32 and the constant learning rate of 1e-5 and 5e-5, respectively. The overall retention rate, most important neuron rate of safe and over-safety patches are set to 30%, 3% and 2%, respectively. As for the hyper-parameter $\alpha$ and $\beta$ in Equation 8, they are set to 1 and 1. In Table 3, we report the average harmful rate on Beavertails, harmful refusal rate on XSTest and performance score on MT-Bench over the three time steps. And in Figure 9, the fine-grained performance changes over the continual alignment process are depicted.
Appendix H Case Study
H.1 Case Study for Safety Enhancement
@@
H.2 Case Study for Over-Safety Mitigation
@@
H.3 Case Study for Utility Preservation (Increasing)
@@
Firstly, the primary reason SafePatching enhances model performance on MT-bench is that it effectively reduces the issue of over-safety in the backbone model, allowing it to directly address the user’s questions. As illustrated in Figure 10, when tasked with named entity recognition (NER), the original model, due to over-safety concerns, would first critique the phrasing of the user’s query (the input sentence of NER). This response is irrelevant to the task at hand, as highlighted by GPT-4o. Our SafePatching, however, successfully addresses this issue, directly meeting the user’s harmless request, and is thus more favored by GPT-4o.
@@
Figure 10: Case study of the generated responses from the orignal backbone model and that from our SafePatching. We also provide the rating and rational from the GPT-4o evaluator.